An investment broker deposits $2000 into an account that earns 4.4% annual interest compounded twice a year. What is the value of the investment after 4 years?
Answer by rfer(16322)

a=2000(1+.044/2)^4*2
a=2000(1.022)^8
a=2000*1.19
a=$2380.33